# Gapless symmetry protected topological phases and generalized deconfined critical points from gauging a finite subgroup

### v1

@@ IV.4 SET phase
Having discussed the SPT phase, we now briefly touch upon the emergent $\tilde{U}(1)$-SET phase (Fig. 4(b)).
@@ IV.5 Generalized DQCP
As we have discussed above, even though $\pi$-vortices are suppressed once the zero-flux condition is enforced, the proliferation of topological phase winding excitations drives the SPT phase to the SET phase. Since the original superfluid-insulator transition is continuous and only a finite subgroup is gauged, we expect this inherited transition to be continuous as well [33]. Thus, we obtain a generalized DQCP from the gapless SPT phase with preserved ^$\hat{\mathbb{Z}}_{2}$ 1-form symmetry and broken $\tilde{U}(1)$ to a SET phase with $\tilde{U}(1)$ (see Fig. 4(b)).
As a result of partial gauging of a finite group, much of the information encoded in the order/disorder parameters can be directly read from the original superfluid-insulator transition. The critical point can be determined by the change in the scaling laws of different order/disorder parameters. The symmetry breaking of higher-form symmetries at critical points has been investigated in the recent literature [62, 63, 64]. If all symmetries involved including the higher-form symmetry are preserved at the critical point in some systems, we may get a 2-D analogue of the intrinsically gapless SPT in 1-D by invoking similar anomaly arguments.
@@
In terms of entanglement entropy $S_{R}$ of a smooth simply connected region $R$ without corners, it is known that, other than the leading perimeter law term $S_{A}\propto l\equiv|\partial R|$, there is a logarithmic subleading correction in a SSB phase of a continuous symmetry [66] and a topological subleading correction in a topologically ordered phase [67, 68]. At the critical point, $S_{A}$ takes the form of $S_{A}=\alpha l-\beta$ with $\beta=\beta_{XY}+\beta_{\mathbb{Z}_{2}}$. Here $\beta_{XY}$ comes from the ordinary $XY$ transition, i.e. SSB of $U(1)$, and $\beta_{\mathbb{Z}_{2}}=\ln 2$ is the topological entanglement entropy of the $\mathbb{Z}_{2}$ topologically ordered phase [50].
@@ IV.6 Perturbations
To study the edge physics, the authors in Ref. [27] numerically demonstrated the robustness of the topological edge modes in the SPT/Higgs phase. On the other hand, the robustness of $\mathbb{Z}_{2}$ topological order with open boundaries under perturbation is also numerically investigated in Ref. [69]. Based on their results, the robustness depends both on the boundary type (rough or smooth) and the perturbation type ($\sigma^{x}$ or $\sigma^{z}$). For rough boundaries (see Fig. 5(a)) and perturbations of the form $\sigma^{x}$, the topological order is robust. It is natural to expect resilience in the edge physics.
@@ V Conclusions
In this work, we investigated the emergent $\mathbb{Z}_{2}$-gauged matter theory of the 1-D and 2-D Bose-Hubbard model coupled to Ising degrees of freedom. We analyzed the inherited phase diagram from that of the ungauged superfluid-insulator version. In 1-D, we identified the superfluid phase to be an intrinsically gapless SPT phase protected by $W\times U(1)$, $W$ being the Ising spin reflection symmetry. In the low energy theory, $W$ can be viewed as the dual ^$\hat{\mathbb{Z}}_{2}$ symmetry. We discussed the effective action which includes a mixed anomaly term between $W$ and the quotient $\tilde{U}(1)$ symmetry $\omega$, and a topological term $\alpha$ dictating the edge degrees of freedom if there is an open boundary. The ’t Hooft anomaly in $\alpha$ is matched by that in $\omega$. We argued that the gapless SPT phase is described by a $\mathbb{Z}_{2}$-gauged compact boson CFT, which is also supported by DMRG computations. In 2-D, we focused on the zero-flux limit and concluded, by adapting the “Higgs = SPT” argument, that the superfluid phase is also a gapless SPT protected by higher form symmetries whose gaplessness comes from the Goldstone modes due to the SSB of the quotient $\tilde{U}(1)$. We studied the excitations, especially the Goldstone modes whose winding number is related to the charge of the $\tilde{U}(1)$ domain wall under the dual ^$\hat{\mathbb{Z}}_{2}$ 1-form symmetry $W$, which is a direct manifestation of the mixed anomaly between the two symmetries. The other phase is the insulating phase corresponding to the SSB of the $W$ with $\tilde{U}(1)$ preserved, i.e. $\tilde{U}(1)$-enriched $\mathbb{Z}_{2}$ topological order. Then we analyzed the transition between the gapless SPT/Higgs phase and the $\tilde{U}(1)$-SET phase, which is a generalized DQCP. The robustness of the gapless SPT/Higgs phase, the SET phase, and the generalized DQCP between them toward perturbations that explicitly break the ^$\hat{\mathbb{Z}}_{2}$ 1-form symmetry is discussed. Possible experimental realizations using quantum simulations with cold atoms are also proposed.
@@
As we have mentioned earlier on, our analysis generalizes easily to fermionic systems. In particular, the 1-D intrinsically gapless SPT phase works for a free fermion gas or a Luttinger liquid and is generalizable to generic critical point or a Fermi liquid in higher dimensions. Generically, introduction of a weakly fluctuating gauge field may destabilize the system and drive the system to other phases, such as superconductors. It would be interesting to construct such a stable intrinsically gapless fermionic SPT phase. A good starting point may be exactly solvable models of free lattice fermions coupled to Ising spins on the link. When it is not analytically solvable, numerical simulations using, e.g., the determinant quantum Monte Carlo method similar to Refs. [94, 95] can give us more valuable insights. We hope our work can stimulate more endeavors along these directions.
@@ Appendix A Emergent $\mathbb{Z}_{2}$-gauged $q$-state clock model
For the Ising model, i.e. when $q=2$, gauging the $\mathbb{Z}_{2}$ is equivalent to a Kramers-Wannier transformation from the original Ising model to the dual Ising model with a dual 0-form ^$\hat{\mathbb{Z}}_{2}$ symmetry. The minimal non-trivial case corresponds to $q=4$, which is also discussed in Refs. [34, 33]. The $\mathbb{Z}_{2}$-gauged 4-state clock model has global symmetry ~$\tilde{\mathbb{Z}}_{2}\times\hat{\mathbb{Z}}_{2}$ with a mixed anomaly between the two symmetries characterized by the non-trivial extension class in $H^{2}(\tilde{\mathbb{Z}}_{2},\hat{\mathbb{Z}}_{2})=\mathbb{Z}_{2}$. A concrete lattice model can be written down as follows,
@@
In the case of $h\gg 1$, $S_{j}$ will be polarized, which implies that $\langle\tau_{m-1/2}^{x}\tau_{n+1/2}^{x}\rangle=\prod_{m\leq j\leq n}S_{j}^{2}\neq 0$, i.e. there is long range order in $\tau^{x}$, leading to SSB of $W$ while $V$ is preserved. On the other hand, when $h\ll 1$, there will be SSB in $V$ but with $W$ preserved. The SSB in $V$ directly inherits from the SSB of the ungauged clock model, since the $\mathbb{Z}_{2}$ gauging corresponds to summing over twisted boundary conditions, which does not change the long-range correlation of the order parameter [34, 33]. Due to the SSB of $V$, which is simply ~$\tilde{\mathbb{Z}}_{2}$ in the bulk, the ground state is 2-fold degenerate $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$ with $V|\psi_{1}\rangle=|\psi_{2}\rangle$.
@@
The analogue of the intrinsically gapless SPT of the $U(1)$ case shows up when $q\geq 5$. It is known that the 1-D quantum clock model without gauging has two critical points, both are of BKT type and dual to each other, and there is a critical phase with emergent $U(1)$ symmetry in between the two critical points [96]. The minimal non-trivial case with mixed anomaly after gauging is the 8-state clock model. After gauging, the emergent symmetry is ~$\tilde{\mathbb{Z}}_{4}\times\hat{\mathbb{Z}}_{2}$, where there is a mixed anomaly characterized by the non-trivial extension class in $H^{2}(\tilde{\mathbb{Z}}_{4},\hat{\mathbb{Z}}_{2})=\mathbb{Z}_{2}$.
Fig. 6 shows the schematic phase diagram with the corresponding ground state degeneracies under both PBC and OBC. In the large $h$ limit, the dual ^$\hat{\mathbb{Z}}_{2}$ symmetry is spontaneously broken, which is labeled as the ^$\hat{\mathbb{Z}}_{2}$ SSB phase. In the small $h$ limit, we have the ~$\tilde{\mathbb{Z}}_{4}$ SSB phase. In the intermediate coupling regime ($h_{1}<h<h_{2}$), the system is in a critical phase, where both of the two symmetries are preserved. There is an emergent symmetry $U(1)\times U(1)\supset\tilde{\mathbb{Z}}_{4}\times\tilde{\mathbb{Z}}_{2}$ and the critical phase is described by the ($\mathbb{Z}_{2}$-gauged) compact boson CFT. Indeed, this phase is an analogue of the intrinsically gapless SPT phase we discussed in the main text.
@@
where $\gamma$ is a loop running along the bonds of the lattice. Notice that in order for there to be an emergent mixed anomaly between $V$ and $W$, the group extension of $V$ by $W$, classified by $H^{2}(\tilde{\mathbb{Z}}_{q/2},\hat{\mathbb{Z}}_{2})=\mathbb{Z}_{\text{gcd}(q/2,2)}$ has to be non-trivial, meaning that $q/2$ has to be even, i.e. $q$ is an integer multiple of 4. Similar to the 1-D case, due to the mixed anomaly between $V$ and $W$, $W$ is SSB while $V$ is preserved when $h\gg 1$. This is the $\mathbb{Z}_{2}$ topological order enriched by ~$\tilde{\mathbb{Z}}_{q/2}$, classified by $H^{2}(\tilde{\mathbb{Z}}_{q/2},\mathbb{Z}_{2}\times\mathbb{Z}_{2})=\mathbb{Z}_{\text{gcd}(q/2,2)}\times\mathbb{Z}_{\text{gcd}(q/2,2)}$.
@@ Appendix B Elimination of gauge constraints
After gauging, i.e. implementing the gauge condition $\prod_{e,i\in\partial e}\sigma^{x}_{e}=(-1)^{n_{i}}$, the onsite boson states are divided into the boson parity even sector and the parity odd sector, depending on the sign of the star operator $\prod_{e,i\in\partial e}\sigma^{x}_{e}$. Therefore, the new onsite boson basis can be denoted as $|1,\tilde{n}\rangle$ and $|-1,\tilde{n}\rangle$, where the first number labels the onsite boson parity, determined by $\prod_{e,i\in\partial e}\sigma^{x}_{e}$, and the second number labels the new local boson states with $\tilde{n}=0,1,2,...$ in the corresponding sector. Notice that we have a one-to-one correspondence between the new basis and the original ungauged basis, given by $|1,\tilde{n}\rangle\leftrightarrow|2\tilde{n}\rangle$ and $|-1,\tilde{n}\rangle\leftrightarrow|2\tilde{n}+1\rangle$. Therefore, no degrees of freedom are lost, as it should be. Expressed in the new basis, the boson number operator becomes
@@
Note that the quantity $\hat{\tilde{N}}=\sum_{i}\hat{\tilde{n}}_{i}$ itself is not conserved. Instead, the term $(1-\prod_{e,i\in\partial e}\sigma^{x}_{e})/2$ contributes a fractional charge 1/2 to $\hat{\tilde{N}}$. By invoking the “electromagnetic duality” for the $\mathbb{Z}_{2}$ gauge field: $\sigma^{x}\leftrightarrow\sigma^{z}$, we can regard the zero-flux condition $\prod_{e\in\partial p}\sigma_{e}^{z}=1$ as the new “Gauss law” for the “gauge field” $\sigma^{x}$ while $\prod_{e,i\in\partial e}\sigma^{x}_{e}$ as the new “magnetic” flux operator. Thus, the new magnetic flux carries a fractional charge under $\hat{\tilde{N}}$, a manifestation of the mixed anomaly. Since the anomalous system may be viewed as a boundary of a SPT phase in the (3+1)-D bulk, as mentioned in Ref. [33], it is natural to envision a discrete realization of the bulk by decorating the magnetic monopoles of the ^$\hat{\mathbb{Z}}_{2}$ gauge field with unit charges under $\tilde{U}(1)$, which is an analogue of the continuum construction in Ref. [98].
@@ Appendix C Grand canonical ensemble in 1-D
In Sec. III, we used canonical enemble (CE) in 1-D for simplicity since the particle number conservation cannot be violated due to the Mermin-Wagner theorem. In this appendix, we support this statement by presenting some DMRG results for the $\mathbb{Z}_{2}$-gauged Bose-Hubbard model in the context of grand canonical ensemble (GCE) where the total particle number can vary. GCE is more general and it allows energy levels consisting of both even and odd parity states (see Fig. 7(a)). The existence of states with different parities is a consequence of the boundaries when using OBC, where the parity operator $P=\sigma_{1/2}\sigma_{L+1/2}$ can still take values ±$\pm 1$. However, GCE is computationally more challenging in DMRG since particle number is not fixed. On the other hand, we observed numerically that as long as the chemical potential $\mu$ is carefully tuned to ensure unit filling for the GCE ground state, both the ground state and the next excited state have no fluctuations in total boson number, i.e. they have fixed boson number. A nice consequence of this observation is that the GCE low energy states can now be related to those of the CE, which renders the numerical calculation easier and more tractable for larger system sizes. We can still show that the superfluid phase is a gapless SPT phase with double degeneracy.
@@
where $\hat{N}$ is the total boson number operator. As mentioned previously, the chemical potential $\mu$ can be tuned to achieve a ground state with unit filling and fixed boson number, i.e. $\langle\hat{N}\rangle=L$ in the ground state. Here $L$ is the number of boson sites which for simplicity is taken to be even. Notice, however, that the proper $\mu$ has a strong size dependence. To carry out a finite size scaling of the gap, we extract the thermodynamical information by bounding the gap as follows.
For a typical set of parameters both the gap in the parity even sector $\Delta_{\text{GCE}}^{P=+1}$ and the gap in the parity odd sector $\Delta_{\text{GCE}}^{P=-1}$ are much larger than the true gap $\Delta_{\text{GCE}}$ which is between a parity even state and a parity odd state (Fig. 7(a)). Note that the ground state is exactly doubly degenerate as in the CE. The GCE ground state energy is given by,
@@
Figure 7: (a) Schematic diagram for the energy levels in the grand canonical ensemble, where parity even and parity odd states coexist. $\Delta_{\text{GCE}}^{P=\pm 1}$ is the gap in the parity even/odd sector and $\Delta_{\mathrm{GCE}}$ is the true gap above the doubly degenerate ground state. (b) The upper bound $\mu_{U}$ and the lower bound $\mu_{L}$ for the chemical potential in order to have unit filling at different system sizes. The inset shows the difference between the two bounds. The parameters used are $t=0.5,U=1.0$.

### v2

@@ IV.4 SET phase
Having discussed the SPT phase, we now briefly touch upon the emergent $\tilde{U}(1)$-SET phase [Fig. 4(b)].
@@ IV.5 Generalized DQCP
As we have discussed above, even though $\pi$-vortices are suppressed once the zero-flux condition is enforced, the proliferation of topological phase winding excitations drives the SPT phase to the SET phase. Since the original superfluid-insulator transition is continuous and only a finite subgroup is gauged, we expect this inherited transition to be continuous as well [33]. Thus, we obtain a generalized DQCP from the gapless SPT phase with preserved ^$\hat{\mathbb{Z}}_{2}$ 1-form symmetry and broken $\tilde{U}(1)$ to a SET phase with $\tilde{U}(1)$ [see Fig. 4(b)].
As a result of partial gauging of a finite group, much of the information encoded in the order/disorder parameters can be directly read from the original superfluid-insulator transition. The critical point can be determined by the change in the scaling laws of different order/disorder parameters. The symmetry breaking of higher-form symmetries at critical points has been investigated in the recent literature [66, 67, 68]. If all symmetries involved including the higher-form symmetry are preserved at the critical point in some systems, we may get a 2-D analog of the intrinsically gapless SPT in 1-D by invoking similar anomaly arguments.
@@
In terms of entanglement entropy $S_{R}$ of a smooth simply connected region $R$ without corners, it is known that, other than the leading perimeter law term $S_{A}\propto l\equiv|\partial R|$, there is a logarithmic subleading correction in a SSB phase of a continuous symmetry [70] and a topological subleading correction in a topologically ordered phase [71, 72]. At the critical point, $S_{A}$ takes the form of $S_{A}=\alpha l-\beta$ with $\beta=\beta_{XY}+\beta_{\mathbb{Z}_{2}}$. Here $\beta_{XY}$ comes from the ordinary $XY$ transition, i.e., SSB of $U(1)$, and $\beta_{\mathbb{Z}_{2}}=\ln 2$ is the topological entanglement entropy of the $\mathbb{Z}_{2}$ topologically ordered phase [54].
@@ IV.6 Perturbations
To study the edge physics, the authors in Ref. [27] numerically demonstrated the robustness of the topological edge modes in the SPT/Higgs phase. On the other hand, the robustness of $\mathbb{Z}_{2}$ topological order with open boundaries under perturbation is also numerically investigated in Ref. [73]. Based on their results, the robustness depends both on the boundary type (rough or smooth) and the perturbation type ($\sigma^{x}$ or $\sigma^{z}$). For rough boundaries [see Fig. 5(a)] and perturbations of the form $\sigma^{x}$, the topological order is robust. It is natural to expect resilience in the edge physics.
@@ V Conclusions
In this work, we investigated the emergent $\mathbb{Z}_{2}$-gauged matter theory of the 1-D and 2-D Bose-Hubbard model coupled to Ising degrees of freedom. We analyzed the inherited phase diagram from that of the ungauged superfluid-insulator version. In 1-D, we identified the superfluid phase to be an intrinsically gapless SPT phase protected by $W\times U(1)$, $W$ being the Ising spin reflection symmetry. In the low energy theory, $W$ can be viewed as the dual ^$\hat{\mathbb{Z}}_{2}$ symmetry. We discussed the effective action which includes a mixed anomaly term between $W$ and the quotient $\tilde{U}(1)$ symmetry $\omega$, and a topological term $\alpha$ dictating the edge degrees of freedom if there is an open boundary. The ’t Hooft anomaly in $\alpha$ is matched by that in $\omega$. We argued that the gapless SPT phase is described by a $\mathbb{Z}_{2}$-gauged compact boson CFT, which is also supported by DMRG computations. In 2-D, we focused on the zero-flux limit and concluded, by adapting the “Higgs = SPT” argument, that the superfluid phase is also a gapless SPT protected by higher form symmetries whose gaplessness comes from the Goldstone modes due to the SSB of the quotient $\tilde{U}(1)$. We studied the excitations, especially the Goldstone modes whose winding number is related to the charge of the $\tilde{U}(1)$ domain wall under the dual ^$\hat{\mathbb{Z}}_{2}$ 1-form symmetry $W$, which is a direct manifestation of the mixed anomaly between the two symmetries. The other phase is the insulating phase corresponding to the SSB of the $W$ with $\tilde{U}(1)$ preserved, i.e., $\tilde{U}(1)$-enriched $\mathbb{Z}_{2}$ topological order. Then we analyzed the transition between the gapless SPT/Higgs phase and the $\tilde{U}(1)$-SET phase, which is a generalized DQCP. The robustness of the gapless SPT/Higgs phase, the SET phase, and the generalized DQCP between them toward perturbations that explicitly break the ^$\hat{\mathbb{Z}}_{2}$ 1-form symmetry is discussed. Possible experimental realizations using quantum simulations with cold atoms are also proposed.
@@
As we have mentioned earlier on, our analysis generalizes easily to fermionic systems. In particular, the 1-D intrinsically gapless SPT phase works for a free fermion gas or a Luttinger liquid [49] and is generalizable to generic critical point or a Fermi liquid in higher dimensions. Generically, introduction of a weakly fluctuating gauge field may destabilize the system and drive the system to other phases, such as superconductors. It would be interesting to construct such a stable intrinsically gapless fermionic SPT phase. A good starting point may be exactly solvable models of free lattice fermions coupled to Ising spins on the link. When it is not analytically solvable, numerical simulations using, e.g., the determinant quantum Monte Carlo method similar to Refs. [98, 99] can give us more valuable insights. We hope our work can stimulate more endeavors along these directions.
@@ Appendix A Emergent $\mathbb{Z}_{2}$-gauged $q$-state clock model
For the Ising model, i.e., when $q=2$, gauging the $\mathbb{Z}_{2}$ is equivalent to a Kramers-Wannier transformation from the original Ising model to the dual Ising model with a dual 0-form ^$\hat{\mathbb{Z}}_{2}$ symmetry. The minimal nontrivial case corresponds to $q=4$, which is also discussed in Refs. [34, 33]. The $\mathbb{Z}_{2}$-gauged 4-state clock model has global symmetry ~$\tilde{\mathbb{Z}}_{2}\times\hat{\mathbb{Z}}_{2}$ with a mixed anomaly between the two symmetries characterized by the non-trivial extension class in $H^{2}(\tilde{\mathbb{Z}}_{2},\hat{\mathbb{Z}}_{2})=\mathbb{Z}_{2}$. A concrete lattice model can be written down as follows,
@@
In the case of $h\gg 1$, $S_{j}$ will be polarized, which implies that $\langle\tau_{m-1/2}^{x}\tau_{n+1/2}^{x}\rangle=\prod_{m\leq j\leq n}S_{j}^{2}\neq 0$, i.e., there is long range order in $\tau^{x}$, leading to SSB of $W$ while $V$ is preserved. On the other hand, when $h\ll 1$, there will be SSB in $V$ but with $W$ preserved. The SSB in $V$ directly inherits from the SSB of the ungauged clock model, since the $\mathbb{Z}_{2}$ gauging corresponds to summing over twisted boundary conditions, which does not change the long-range correlation of the order parameter [34, 33]. Due to the SSB of $V$, which is simply ~$\tilde{\mathbb{Z}}_{2}$ in the bulk, the ground state is 2-fold degenerate $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$ with $V|\psi_{1}\rangle=|\psi_{2}\rangle$.
@@
The analog of the intrinsically gapless SPT of the $U(1)$ case shows up when $q\geq 5$. It is known that the 1-D quantum clock model without gauging has two critical points, both are of BKT type and dual to each other, and there is a critical phase with emergent $U(1)$ symmetry in between the two critical points [100]. The minimal non-trivial case with mixed anomaly after gauging is the 8-state clock model. After gauging, the emergent symmetry is ~$\tilde{\mathbb{Z}}_{4}\times\hat{\mathbb{Z}}_{2}$, where there is a mixed anomaly characterized by the non-trivial extension class in $H^{2}(\tilde{\mathbb{Z}}_{4},\hat{\mathbb{Z}}_{2})=\mathbb{Z}_{2}$.
Fig. 6 shows the schematic phase diagram with the corresponding ground state degeneracies under both PBC and OBC. In the large $h$ limit, the dual ^$\hat{\mathbb{Z}}_{2}$ symmetry is spontaneously broken, which is labeled as the ^$\hat{\mathbb{Z}}_{2}$ SSB phase. In the small $h$ limit, we have the ~$\tilde{\mathbb{Z}}_{4}$ SSB phase. In the intermediate coupling regime ($h_{1}<h<h_{2}$), the system is in a critical phase, where both of the two symmetries are preserved. There is an emergent symmetry $U(1)\times U(1)\supset\tilde{\mathbb{Z}}_{4}\times\tilde{\mathbb{Z}}_{2}$ and the critical phase is described by the ($\mathbb{Z}_{2}$-gauged) compact boson CFT. Indeed, this phase is an analog of the intrinsically gapless SPT phase we discussed in the main text.
@@
where $\gamma$ is a loop running along the bonds of the lattice. Notice that in order for there to be an emergent mixed anomaly between $V$ and $W$, the group extension of $V$ by $W$, classified by $H^{2}(\tilde{\mathbb{Z}}_{q/2},\hat{\mathbb{Z}}_{2})=\mathbb{Z}_{\text{gcd}(q/2,2)}$ has to be non-trivial, meaning that $q/2$ has to be even, i.e., $q$ is an integer multiple of 4. Similar to the 1-D case, due to the mixed anomaly between $V$ and $W$, $W$ is SSB while $V$ is preserved when $h\gg 1$. This is the $\mathbb{Z}_{2}$ topological order enriched by ~$\tilde{\mathbb{Z}}_{q/2}$, classified by $H^{2}(\tilde{\mathbb{Z}}_{q/2},\mathbb{Z}_{2}\times\mathbb{Z}_{2})=\mathbb{Z}_{\text{gcd}(q/2,2)}\times\mathbb{Z}_{\text{gcd}(q/2,2)}$.
@@ Appendix B Elimination of gauge constraints
After gauging, i.e., implementing the gauge condition $\prod_{e,i\in\partial e}\sigma^{x}_{e}=(-1)^{n_{i}}$, the onsite boson states are divided into the boson parity even sector and the parity odd sector, depending on the sign of the star operator $\prod_{e,i\in\partial e}\sigma^{x}_{e}$. Therefore, the new onsite boson basis can be denoted as $|1,\tilde{n}\rangle$ and $|-1,\tilde{n}\rangle$, where the first number labels the onsite boson parity, determined by $\prod_{e,i\in\partial e}\sigma^{x}_{e}$, and the second number labels the new local boson states with $\tilde{n}=0,1,2,...$ in the corresponding sector. Notice that we have a one-to-one correspondence between the new basis and the original ungauged basis, given by $|1,\tilde{n}\rangle\leftrightarrow|2\tilde{n}\rangle$ and $|-1,\tilde{n}\rangle\leftrightarrow|2\tilde{n}+1\rangle$. Therefore, no degrees of freedom are lost, as it should be. Expressed in the new basis, the boson number operator becomes
@@
Note that the quantity $\hat{\tilde{N}}=\sum_{i}\hat{\tilde{n}}_{i}$ itself is not conserved. Instead, the term $(1-\prod_{e,i\in\partial e}\sigma^{x}_{e})/2$ contributes a fractional charge 1/2 to $\hat{\tilde{N}}$. By invoking the “electromagnetic duality” for the $\mathbb{Z}_{2}$ gauge field: $\sigma^{x}\leftrightarrow\sigma^{z}$, we can regard the zero-flux condition $\prod_{e\in\partial p}\sigma_{e}^{z}=1$ as the new “Gauss law” for the “gauge field” $\sigma^{x}$ while $\prod_{e,i\in\partial e}\sigma^{x}_{e}$ as the new “magnetic” flux operator. Thus, the new magnetic flux carries a fractional charge under $\hat{\tilde{N}}$, a manifestation of the mixed anomaly. Since the anomalous system may be viewed as a boundary of an SPT phase in the (3+1)-D bulk, as mentioned in Ref. [33], it is natural to envision a discrete realization of the bulk by decorating the magnetic monopoles of the ^$\hat{\mathbb{Z}}_{2}$ gauge field with unit charges under $\tilde{U}(1)$, which is an analog of the continuum construction in Ref. [102].
@@ Appendix C Grand canonical ensemble in 1-D
In Sec. III, we used canonical enemble (CE) in 1-D for simplicity since the particle number conservation cannot be violated due to the Mermin-Wagner theorem. In this appendix, we support this statement by presenting some DMRG results for the $\mathbb{Z}_{2}$-gauged Bose-Hubbard model in the context of grand canonical ensemble (GCE) where the total particle number can vary. GCE is more general and it allows energy levels consisting of both even and odd parity states [see Fig. 7(a)]. The existence of states with different parities is a consequence of the boundaries when using OBC, where the parity operator $P=\sigma_{1/2}\sigma_{L+1/2}$ can still take values ±$\pm 1$. However, GCE is computationally more challenging in DMRG since particle number is not fixed. On the other hand, we observed numerically that as long as the chemical potential $\mu$ is carefully tuned to ensure unit filling for the GCE ground state, both the ground state and the next excited state have no fluctuations in total boson number, i.e., they have fixed boson number. A nice consequence of this observation is that the GCE low energy states can now be related to those of the CE, which renders the numerical calculation easier and more tractable for larger system sizes. We can still show that the superfluid phase is a gapless SPT phase with double degeneracy.
@@
where $\hat{N}$ is the total boson number operator. As mentioned previously, the chemical potential $\mu$ can be tuned to achieve a ground state with unit filling and fixed boson number, i.e., $\langle\hat{N}\rangle=L$ in the ground state. Here $L$ is the number of boson sites which for simplicity is taken to be even. Notice, however, that the proper $\mu$ has a strong size dependence. To carry out a finite size scaling of the gap, we extract the thermodynamical information by bounding the gap as follows.
For a typical set of parameters both the gap in the parity even sector $\Delta_{\text{GCE}}^{P=+1}$ and the gap in the parity odd sector $\Delta_{\text{GCE}}^{P=-1}$ are much larger than the true gap $\Delta_{\text{GCE}}$ which is between a parity even state and a parity odd state [Fig. 7(a)]. Note that the ground state is exactly doubly degenerate as in the CE. The GCE ground state energy is given by,
@@
Figure 7: (a) Schematic diagram for the energy levels in the grand canonical ensemble, where parity even and parity odd states coexist. $\Delta_{\text{GCE}}^{P=\pm 1}$ is the gap in the parity even/odd sector and $\Delta_{\mathrm{GCE}}$ is the true gap above the doubly degenerate ground state. (b) The upper bound $\mu_{U}$ and the lower bound $\mu_{L}$ for the chemical potential in order to have unit filling at different system sizes. The inset shows the difference between the two bounds. The parameters used are $t=0.5$ and $U=1.0$.